# PALM: A Dataset and Baseline for Learning Multi-subject Hand Prior

### v2

@@ 4.2 PALM-Net
Physically based representations: Inspired by [wang2024intrinsicavatar, chen2022gdna], PALM-Net (Figure 4c) decomposes the hand representation into a shape network $f_{g}(\cdot)$, a radiance field network $f_{rf}(\cdot)$, and a material network $f_{m}(\cdot)$. We model the canonical hand geometry with an implicit function:
@@
where $\mathbf{x}_{c}\in\rm I\!R^{3}$ denotes a 3D point in the canonical space, $\theta,\beta$ are MANO parameter, and $\phi\in\rm I\!R^{d_{s}}$ captures subject-specific geometry latent code of dimension $d_{s}$. $f_{g}(\cdot)$ outputs the opacity value, $\sigma_{t}(\mathbf{x}_{c})\in\rm I\!R$, as well as geometry features $z(\mathbf{x}_{c})\in\rm I\!R^{d_{s}}$ for the point $\mathbf{x}_{c}$. Following [yariv2021volume], the opacity is obtained by converting the Signed Distance Function (SDF) values via the cumulative distribution function of the scaled Laplace distribution, $\Gamma_{\alpha_{1},\alpha_{2}}(s)$, where $\alpha_{1},\alpha_{2}>0$ are optimizable parameters. For details, we refer the reader to [yariv2021volume].
@@
Physically based rendering: For physically based rendering, we follow closely [wang2024intrinsicavatar] and compute the radiance scattered by the volume along a certain camera ray ($\mathbf{o,d}$) using the quadrature approximation as,
@@
where $(\bar{t}^{(1)},\bar{t}^{(2)},...,\bar{t}^{(M)})$ are the importance sampling offsets from the PDF estimated by radiance field samples in Equation 2; $M$ denotes the number of samples used to approximate the integrals along the ray; $\bar{\mathbf{d}}^{(i)}$ is the incoming light direction at sampling offset $\bar{t}^{(i)}$ sampled from the distribution, $pdf(\cdot)$ (uniform distribution over the unit sphere); $BRDF(\cdot)$ denotes the simplified version of Disney BRDF [Burley2012SIGGRAPH]. The term $L_{i}(\mathbf{x},\bar{\mathbf{d}})$ is the incoming radiance towards point $\mathbf{x}$ along direction $\bar{\mathbf{d}}$ and can be computed as the weighted sum of output radiance $C_{rf}(\mathbf{x},\bar{\mathbf{d}})$ (Equation 2) and radiance emitted from an environment map $\text{Env}(\bar{\mathbf{d}})$:
@@
where ${t_{n}^{\prime}},{t_{f}^{\prime}}$ are near and far points of integration for secondary rays. The environment map is approximated by a set of Spherical Gaussians denoted by $\mathcal{SG}_{1},\mathcal{SG}_{2},...,\mathcal{SG}_{G}$. We refer the reader to [wang2024intrinsicavatar] for derivation of the above equations.
@@
Note that the geometry information is shared by PBR and radiance field. We encourage valid SDFs with the eikonal loss $\mathcal{L}_{\text{eikonal}}$ [gropp2020igr], which enforces the gradient at each point to have a unit norm. To encourage smooth hand surfaces, we apply a Laplacian loss $\mathcal{L}_{\text{LAP}}$ on sampled points around the hand (see SupMat). To encourage latent codes to be close to zeros, we penalize a MSE loss $\mathcal{L}_{\text{latent}}$ on both appearance and shape code. To avoid foreground model explaining background pixels, we also supervise the networks with a segmentation loss
@@
where $\mathcal{S}(\mathbf{r})\in\rm I\!R$ represents the probability of a pixel being the foreground and $\text{BCE}(\cdot,\cdot)$ is the binary cross entropy loss to the ground-truth hand segmentation mask $\hat{\mathcal{S}}(\mathbf{r})$ rendered from the hand scans. Finally, to capture high-frequency details, we render image patches and compare them with the ground-truth using the perceptual similarity loss $\mathcal{L}_{\text{LPIPS}}$ [zhang2018perceptual]. The total loss $\mathcal{L}$ is defined as
@@ 5.1 Datasets
InterHand2.6M: To evaluate the performance on real images, we use images from InterHand2.6M [interhand] for personalization. The dataset consists of accurate 3D hand poses of subjects in predefined poses as well as high-resolution RGB images. We evaluate on right-hand only sequences using the test split of the dataset. To reliably measure performance, we randomly select two views for each sequence, resulting in 12 sequences. We uniformly sample 20 images for each sequence for the evaluation and use the first frame of each sequence for training a personalized model.
HARP relit: Since there is no real dataset to evaluate hand avatar under novel environment and poses, following [karunratanakul2023harp], we render a synthetic dataset using Blender. In particular, to create synthetic hand template, for each sequence, we sample new MANO shape parameters and apply a new skin tone using UV textures from DART [gao2022dart]. To animate the hand, we use the hand pose parameters from HARP data release [karunratanakul2023harp]. We use the first frame of each sequence for training a personalized model and use the remaining frames for evaluation. We render the training and evaluation images using different environment maps to evaluate relighting.
In-the-wild images: To show the generalization of our method under real world scenarios, we select in-the-wild images from the internet with diverse lighting conditions and poses. After personalization on these images, we render them with novel environment maps using novel poses from [karunratanakul2023harp] and show qualitative results.
@@
  Handy [potamias2023handy]   7.50   0.69   0.24
  HARP [karunratanakul2023harp]   9.89   0.78   0.16
  UHM [moon2024authentic]   10.08   0.76   0.19
@@
  Handy [potamias2023handy]   12.48   0.76   0.32
  HARP [karunratanakul2023harp]   11.93   0.69   0.37
  UHM [moon2024authentic]   12.30   0.74   0.31
@@ 5.2 Comparison and Analysis
Supervision with 3dMD normals: Traditional multi-view techniques require a massive amount of RGB views for high-fidelity 3D reconstruction [mildenhall2021nerf]. In our capture setup, we use a hybrid approach, combining 3dMD scans and sparse multi-view RGB images for training our prior models. Figure 7 shows a qualitative comparison with and without 3dMD normals. In particular, we train two multi-subject PBR prior models on PALM, one with 3dMD normal supervision and one without. Then we take these two prior models and personalize on an InterHand2.6M image. Figure 7 shows that, 3dMD normals from hand scans are crucial in reducing pepper-like artifacts and it helps to reduce floaters. Table 4 quantitatively compares in this novel pose evaluation.

### v3

@@ 4.2 PALM-Net
Physically based representations: Inspired by [46, 6], PALM-Net (Figure 4c) decomposes the hand representation into a shape network $f_{g}(\cdot)$, a radiance field network $f_{rf}(\cdot)$, and a material network $f_{m}(\cdot)$. We model the canonical hand geometry with an implicit function:
@@
where $\mathbf{x}_{c}\in\rm I\!R^{3}$ denotes a 3D point in the canonical space, $\theta,\beta$ are MANO parameter, and $\phi\in\rm I\!R^{d_{s}}$ captures subject-specific geometry latent code of dimension $d_{s}$. $f_{g}(\cdot)$ outputs the opacity value, $\sigma_{t}(\mathbf{x}_{c})\in\rm I\!R$, as well as geometry features $z(\mathbf{x}_{c})\in\rm I\!R^{d_{s}}$ for the point $\mathbf{x}_{c}$. Following [48], the opacity is obtained by converting the Signed Distance Function (SDF) values via the cumulative distribution function of the scaled Laplace distribution, $\Gamma_{\alpha_{1},\alpha_{2}}(s)$, where $\alpha_{1},\alpha_{2}>0$ are optimizable parameters. For details, we refer the reader to [48].
@@
Physically based rendering: For physically based rendering, we follow closely [46] and compute the radiance scattered by the volume along a certain camera ray ($\mathbf{o,d}$) using the quadrature approximation as,
@@
where $(\bar{t}^{(1)},\bar{t}^{(2)},...,\bar{t}^{(M)})$ are the importance sampling offsets from the PDF estimated by radiance field samples in Equation 2; $M$ denotes the number of samples used to approximate the integrals along the ray; $\bar{\mathbf{d}}^{(i)}$ is the incoming light direction at sampling offset $\bar{t}^{(i)}$ sampled from the distribution, $pdf(\cdot)$ (uniform distribution over the unit sphere); $BRDF(\cdot)$ denotes the simplified version of Disney BRDF [3]. The term $L_{i}(\mathbf{x},\bar{\mathbf{d}})$ is the incoming radiance towards point $\mathbf{x}$ along direction $\bar{\mathbf{d}}$ and can be computed as the weighted sum of output radiance $C_{rf}(\mathbf{x},\bar{\mathbf{d}})$ (Equation 2) and radiance emitted from an environment map $\text{Env}(\bar{\mathbf{d}})$:
@@
where ${t_{n}^{\prime}},{t_{f}^{\prime}}$ are near and far points of integration for secondary rays. The environment map is approximated by a set of Spherical Gaussians denoted by $\mathcal{SG}_{1},\mathcal{SG}_{2},...,\mathcal{SG}_{G}$. We refer the reader to [46] for derivation of the above equations.
@@
Note that the geometry information is shared by PBR and radiance field. We encourage valid SDFs with the eikonal loss $\mathcal{L}_{\text{eikonal}}$ [17], which enforces the gradient at each point to have a unit norm. To encourage smooth hand surfaces, we apply a Laplacian loss $\mathcal{L}_{\text{LAP}}$ on sampled points around the hand (see SupMat). To encourage latent codes to be close to zeros, we penalize a MSE loss $\mathcal{L}_{\text{latent}}$ on both appearance and shape code. To avoid foreground model explaining background pixels, we also supervise the networks with a segmentation loss
@@
where $\mathcal{S}(\mathbf{r})\in\rm I\!R$ represents the probability of a pixel being the foreground and $\text{BCE}(\cdot,\cdot)$ is the binary cross entropy loss to the ground-truth hand segmentation mask $\hat{\mathcal{S}}(\mathbf{r})$ rendered from the hand scans. Finally, to capture high-frequency details, we render image patches and compare them with the ground-truth using the perceptual similarity loss $\mathcal{L}_{\text{LPIPS}}$ [54]. The total loss $\mathcal{L}$ is defined as
@@ 5.1 Datasets
InterHand2.6M: To evaluate the performance on real images, we use images from InterHand2.6M [32] for personalization. The dataset consists of accurate 3D hand poses of subjects in predefined poses as well as high-resolution RGB images. We evaluate on right-hand only sequences using the test split of the dataset. To reliably measure performance, we randomly select two views for each sequence, resulting in 12 sequences. We uniformly sample 20 images for each sequence for the evaluation and use the first frame of each sequence for training a personalized model.
HARP relit: Since there is no real dataset to evaluate hand avatar under novel environment and poses, following [21], we render a synthetic dataset using Blender. In particular, to create synthetic hand template, for each sequence, we sample new MANO shape parameters and apply a new skin tone using UV textures from DART [15]. To animate the hand, we use the hand pose parameters from HARP data release [21]. We use the first frame of each sequence for training a personalized model and use the remaining frames for evaluation. We render the training and evaluation images using different environment maps to evaluate relighting.
In-the-wild images: To show the generalization of our method under real world scenarios, we select in-the-wild images from the internet with diverse lighting conditions and poses. After personalization on these images, we render them with novel environment maps using novel poses from [21] and show qualitative results.
@@
  Handy [35]   7.50   0.69   0.24
  HARP [21]   9.89   0.78   0.16
  UHM [31]   10.08   0.76   0.19
@@
  Handy [35]   12.48   0.76   0.32
  HARP [21]   11.93   0.69   0.37
  UHM [31]   12.30   0.74   0.31
@@ 5.2 Comparison and Analysis
Supervision with 3dMD normals: Traditional multi-view techniques require a massive amount of RGB views for high-fidelity 3D reconstruction [28]. In our capture setup, we use a hybrid approach, combining 3dMD scans and sparse multi-view RGB images for training our prior models. Figure 7 shows a qualitative comparison with and without 3dMD normals. In particular, we train two multi-subject PBR prior models on PALM, one with 3dMD normal supervision and one without. Then we take these two prior models and personalize on an InterHand2.6M image. Figure 7 shows that, 3dMD normals from hand scans are crucial in reducing pepper-like artifacts and it helps to reduce floaters. Table 4 quantitatively compares in this novel pose evaluation.